OpenStudy (anonymous):

15x^2+20/x-6/6x^2+8/x^2-36

OpenStudy (anonymous):

\[\frac{15x^2+20}{x-6}=\frac{6x^2+8}{x^2-36}\] so what are the denominators we have here?

jhonyy9 (jhonyy9):

factoriz x^2 -36 = ... us formula a^2 -b^2 =(a-b)(a+b) in case of numerators factorize out the common factor see what you get on left and right side simplifies what is possible